# BondMatcher: H-Bond Stability Analysis in Molecular Systems

### v1

@@ 5.6 Limitations
A first limitation, which prevented us from analyzing vibrations modes beyond mode $36$ for the Cage and Prism, is that our input database contains a few datasets lacking some critical points corresponding to hydrogen atoms, with a number of local minima of $\rho^{\prime}$ initially smaller than $18$. These are high-frequency modes describing the O--H stretching involving the non-H-bonded H atoms. These modes may be susceptible to effects not explicitly handled in the harmonic oscillator model. Moreover, our model artifically extends the viable range of these motion amplitudes. Thus, the displacements of these H atoms may be overestimated, such that their corresponding minima may sometimes vanish. This might be addressed by re-generating the data on a finer grid (e.g., $512^{3}$), at the expense of larger computation times. Another possibility would be to consider an unbalanced variant of the assignment problem from Sec. 4.2. However, this would result in certain covalent bonds (involving the missing hydrogen atoms) being reported as unstable, which would challenge the overall interpretation of the results.
@@ 6 Conclusion
  • The saddles defining H-bond paths that are unstable under vibrations are also the saddles associated to the lowest persistence (Fig. 8). This indicates that low-persistence saddle pairs are likely to be canceled first by vibrations. However, in principle, path instability can also be observed for persistent saddles (see Fig. 4).
@@
We believe our work opens several research avenues at the intersection between DMT and QTAIM. First, our BondMatcher framework can be used for other studies requiring a robust, per-bond comparative analysis of several systems. For instance, the protocol used in our case studies could be extended to more complex (or more diversified) molecular systems, to get a refined picture of the instability of H-bond paths. Second, our framework could also be used for the comparative analysis of a single molecular system, but studied under various descriptors (e.g., electron density versus bare nuclear potential). Such feature correspondences between descriptors could be instrumental to accelerate several high-throughput chemical analysis tasks, as the electron density can still be computationally expensive to obtain. However, such comparisons would require to revisit our extremum graph matching procedure, in particular to reliably account for distinct extremum counts.

### v2

@@ 5.5 Limitations
A first limitation, which prevented us from analyzing vibrations modes beyond mode $36$ for the Cage and Prism, is that our input database contains a few datasets lacking some critical points corresponding to hydrogen atoms, with a number of local minima of $\rho^{\prime}$ initially smaller than $18$. These are high-frequency modes describing the O--H stretching involving the non-H-bonded H atoms. These modes may be susceptible to effects not explicitly handled in the harmonic oscillator model. Moreover, our model artifically extends the viable range of these motion amplitudes. As a result, the displacements of these H atoms may be overestimated, such that their corresponding minima may sometimes vanish. This might be addressed by re-generating the data on a finer grid (e.g., $512^{3}$), at the expense of larger computation times. Another possibility would be to consider an unbalanced variant of the assignment problem from Sec. 4.2. Nevertheless, this would result in certain covalent bonds (involving the missing hydrogen atoms) being reported as unstable, which would challenge the overall interpretation.
@@ 6 Conclusion
  • The saddles defining H-bond paths that are unstable under vibrations are also the saddles associated to the lowest persistence (Fig. 8). This indicates that low-persistence saddle pairs are likely to be canceled first by vibrations. Still in principle, path instability can also be observed for persistent saddles (see Fig. 4).
@@
We believe our work opens several research avenues at the intersection between DMT and QTAIM. First, our BondMatcher framework can be used for other studies requiring a robust, per-bond comparative analysis of several systems. For instance, the protocol used in our case studies could be extended to more complex molecular systems (e.g., as in [5]), to get a refined picture of the instability of H-bond paths. Second, our framework could also be used for the comparative analysis of a single molecular system, but studied under various descriptors (e.g., electron density versus bare nuclear potential). Such feature correspondences between descriptors could be instrumental to accelerate several high-throughput chemical analysis tasks, as the electron density can still be computationally expensive to obtain. However, such comparisons would require to revisit our extremum graph matching procedure, in particular to reliably account for distinct extremum counts.